# Physics-Informed Neural Networks vs. Physics Models for Non-Invasive Glucose Monitoring: A Comparative Study Under Realistic Synthetic Conditions

### v1

@@ Abstract
Non-invasive glucose monitors often fail outside the lab because existing datasets ignore hardware noise, environmental drift and person-to-person physiology. We introduce the first ultra-realistic near-infrared (NIR) simulator that injects 12-bit ADC quantisation, ±0.1% LED ageing, photodiode dark noise, 15–45°C / 30–90% RH, contact-pressure variation, Fitzpatrick I–VI melanin and diurnal glucose excursions (dawn phenomenon). Using this platform ($\rho_{\text{glucose-NIR}}=0.21$), we benchmark six methods: Enhanced Beer–Lambert (physics-engineered ridge regression), three physics-informed neural networks (PINNs), a selective radiative-transfer PINN, and a shallow DNN. Beer–Lambert achieves 13.6 mg/dL RMSE, 95.8% Clarke-A and 93.8% ±15% accuracy with only 56 parameters and 0.01 ms inference, outperforming the best PINN (14.6 mg/dL) and the SDNN baseline (35.1 mg/dL). Results overturn the assumption that deeper PINNs dominate and supply an open, end-to-end reference stack for rapid prototyping of embedded optical glucose sensors.
@@ 1 Introduction
Non-invasive glucose monitoring remains a long-standing challenge in biomedical sensing, with the potential to transform diabetes management by eliminating painful finger pricks, reducing consumable costs, and enabling real-time glucose tracking. Despite decades of research, no solution has matched the accuracy and reliability of invasive methods in real-world deployment. Near-infrared (NIR) spectroscopy — leveraging glucose’s absorption features in the 850–1150 nm range — remains the most promising modality, but progress has stalled due to the problem’s intrinsic complexity.
@@
Recent efforts have turned to Physics-Informed Neural Networks (PINNs), which embed physical constraints (e.g., Beer–Lambert law or full radiative transport equations) into neural architectures. These models aim to improve generalization by aligning learning with known optics. However, their robustness in harsh, realistic settings — and how they compare to simpler models with strong physics-based features — remains poorly understood.
This paper closes that gap by introducing a high-fidelity synthetic data generator that simulates ultra-realistic hardware, environmental, and physiological variation. Our simulator includes 12-bit ADC quantization, $\pm 0.1\%$ LED drift, contact-pressure noise, Fitzpatrick I–VI melanin, skin thickness from 0.5–4 mm, and diurnal glucose surges (dawn phenomenon). Using this platform, we benchmark six model families:
@@
  • Full RTE PINN (explicit light transport)
@@
Our findings reveal that the Enhanced Beer–Lambert model achieves 13.6 mg/dL RMSE with only 56 parameters and 0.01 ms inference time — outperforming deeper PINNs and SDNNs by a wide margin under realistic correlation ceilings. While our simulator captures a wide spectrum of hardware, environmental, and physiological effects, real-world deployment may still introduce unmodeled complexities — including sensor-skin misalignment, unmeasured confounders (e.g. sweat, hair, dynamic tissue hydration), and population drift over time.
@@
    We develop the first ultra-realistic NIR simulator that integrates hardware noise, environmental drift, and physiological diversity — including dawn effect and melanin variation.
@@
    We perform a rigorous side-by-side evaluation of six models under identical stress-test conditions.
@@
    We show that physics-based feature engineering (Beer–Lambert + PMF) outperforms deeper black-box PINNs, even when all are exposed to the same complex conditions.
@@
To our knowledge, this is the first work to combine Beer–Lambert physics with personalized medical features (PMFs) in a deployable ridge-regression pipeline for non-invasive glucose sensing.
@@ 2 Background and Related Work
2.1 Near-Infrared Glucose-Monitoring Fundamentals
@@
These factors reduce raw glucose–NIR correlation from >0.8>\!0.8 in cuvette studies to $\rho\!\approx\!0.2$ in realistic prototypes, motivating the physics-informed and feature-engineered approaches evaluated in this work.
@@ 2.2 Related Computational Approaches
Deep neural networks (DNN/SDNN). Data-driven models (e.g. Srichan et al. [1]) learn spectral–glucose mappings directly, achieving good results in high-SNR datasets, yet overfit or misgeneralise in realistic noise regimes.
@@
To date, no study rigorously benchmarks physics-engineered linear models against multiple PINN variants under identical, ultra-realistic noise and physiology. The present work fills this gap, showing that carefully engineered Beer–Lambert features coupled with ridge regression can outperform deeper networks in both accuracy and deployability.
@@ 2.3 Physics-Informed Neural Networks
$\mathcal{L}_{\text{Beer\textendash Lambert}}=\frac{1}{N}\sum_{i=1}^{N}\sum_{\lambda}\bigl{|}r_{BL}^{(i,\lambda)}\bigr{|}^{2}.$ (4)
Radiative-transfer (RTE) PINN.
@@
leading to the physics loss
$\mathcal{L}_{\text{RTE}}=\frac{1}{N}\sum_{i=1}^{N}\!\left\|\bm{\Omega}\!\cdot\!\nabla I^{(i)}+\bigl{(}\mu_{a}+\mu_{s}\bigr{)}I^{(i)}-\mu_{s}\!\int\!pI^{(i)}d\bm{\Omega}^{\prime}-S\right\|^{2}.$ (5)
Conservation constraints.
@@
$\mathcal{L}_{\text{cons}}=\frac{1}{N}\sum_{i=1}^{N}\left|\frac{\partial c^{(i)}}{\partial t}+\nabla\!\cdot\!\bigl{(}c^{(i)}\mathbf{v}^{(i)}\bigr{)}\right|^{2},$ (6)
where $\mathbf{v}$ is a glucose–transport velocity field.
@@
PINNs excel when the physics is accurate and complete; otherwise the residual term can mis-guide training. Choosing $\lambda_{\text{phys}}$ is non-trivial: too small and the network ignores physics; too large and it overfits imperfect equations. Moreover, solving Eq. (5) requires dense angular quadrature, inflating both parameter count and training cost.
In the glucose-sensing literature, most reported PINNs include only Beer–Lambert penalties and omit scattering or personalised anatomy, leaving open the question of their advantage under realistic signal-to-noise ratios. Our study quantifies that gap by comparing three PINN variants—including a full RTE PINN—against physics-engineered linear models and an SDNN baseline.
@@ 2.4 Radiative-Transfer Equation Modeling
The Beer–Lambert law assumes a single, straight optical path and no scattering—an assumption violated in biological tissue, where photons scatter tens to hundreds of times before detection. To capture this complexity, the radiative-transfer equation (RTE) models the balance of light energy at every point $\mathbf{r}$, time $t$, and direction $\bm{\Omega}$:
@@
  • - $I(\mathbf{r},\bm{\Omega},t)$ is the radiance — energy per unit area per unit solid angle.
@@
We use the full RTE as a physics constraint in one of our PINN models (3.2.3). This tests whether embedding scattering physics improves generalisation under realistic, noisy optical conditions — and how this compares to empirical compensation in engineered models like EBL.
@@ 3 Methodology
3.1 Ultra-Realistic Synthetic Data Generation Framework
Robust deployment demands that algorithms be stress-tested under the same noise, drift, and physiology a wearable will face in the wild. We therefore build a three-layer simulator—hardware/environment/physiology—and drive realistic glucose trajectories through it. The resulting correlation ceiling ($\rho_{\text{glucose-NIR}}\approx 0.21$) matches values reported for early field prototypes [9, 8].
@@ 3.1.2 Environmental layer: lab–to–field translation
Context: these ranges cover an ICU ward, a tropical street market, and a Colorado ski slope—three environments where non-invasive monitors have previously failed clinical accuracy tests.
@@ 3.1.3 Physiology layer: population and state diversity
3.1.4 Glucose trajectories: meals, dawn surge, exercise
@@
[…]
A stochastic interstitial-lag filter ($\tau=7\!-\!15\;\text{min}$) links plasma to tissue glucose, matching CGM literature.
Outcome: once all layers interact, raw glucose–NIR correlation collapses from 0.82 (cuvette) to 0.21 (prototype), recreating the operating point that has stalled past clinical trials.
@@ 3.2.1 Enhanced Beer–Lambert Model
The 56 features include raw absorbances, selected wavelength contrasts, second-order curvature terms, PMF-weighted components, and direct PMF entries—explicitly reflecting all biologically and physically grounded variables used during deployment.
@@ 3.2.2 Physics-Informed Neural Network Architectures
$\mathcal{L}_{\text{PINN}}=\mathcal{L}_{\text{data}}+\lambda_{\text{physics}}\cdot\mathcal{L}_{\text{Beer\textendash Lambert}},$ (26)
@@
$\mathcal{L}_{\text{Beer\textendash Lambert}}=\frac{1}{N}\sum_{i=1}^{N}\sum_{\lambda}\left|\log\left(\frac{I_{0}}{I_{\lambda}^{(i)}}\right)-\varepsilon(\lambda)\cdot\hat{c}_{\text{glucose}}^{(i)}\cdot l^{(i)}\right|^{2}.$ (27)
This acts as a structural prior, penalizing predictions that violate expected absorbance behavior given optical path length $l$ (we set $l=1$ during training, absorbing relative scale into the learned bias). Setting $l=1$ serves as a normalization strategy rather than a strict physical assertion. Since true optical path lengths vary significantly across individuals due to factors like skin thickness, fat distribution, and hydration, fixing $l$ allows us to preserve the linear form of Beer–Lambert while absorbing inter-subject variability into the learned weights. In practice, the physics loss is softly weighted ($\lambda_{\text{physics}}\ll 1$) so that it acts as a regularizer—nudging the model toward optical consistency without overpowering the data-driven objective, while incorporating known extinction coefficients $\varepsilon(\lambda)$.
To address limitations of vanilla MLPs and improve training dynamics, we augment the PINN with three key architectural innovations:
  • Residual connections, which stabilize gradient flow in deeper layers:
    $\mathbf{h}_{l+1}=\mathbf{h}_{l}+\mathcal{F}(\mathbf{h}_{l},\mathbf{W}_{l}),$ (28)
    where $\mathcal{F}$ is a learned transformation and $\mathbf{W}_{l}$ the corresponding weights.
    This structure allows the network to learn only the residual (i.e., the change or correction) needed between layers, rather than the entire transformation from scratch. During training, if a layer’s transformation $\mathcal{F}$ learns nothing useful, the network still has the identity path $\mathbf{h}_{l+1}\approx\mathbf{h}_{l}$, preserving signal flow and preventing degradation. In practice, this makes deeper models easier to optimize, reduces vanishing gradients, and supports more expressive representations. Conceptually, each layer acts more like a “refinement” of what came before rather than a complete rewrite—crucial when learning subtle physiological correlations from high-dimensional, noisy NIR inputs.
  • Spectral attention, which adaptively reweights wavelengths by learned relevance:
    $\alpha_{\lambda}=\frac{\exp(e_{\lambda})}{\sum_{\lambda^{\prime}}\exp(e_{\lambda^{\prime}})},\quad\text{with}\quad e_{\lambda}=\mathbf{w}_{\text{attn}}^{\top}\cdot A_{\lambda},$ (29)
    where $A_{\lambda}$ is the log absorbance and $\alpha_{\lambda}$ modulates feature emphasis across the NIR band.
    This formulation mimics a soft attention mechanism, where each wavelength is scored by a learned relevance term $e_{\lambda}$, then normalized using a softmax to produce attention weights $\alpha_{\lambda}$. Intuitively, this allows the model to "focus" on the most informative parts of the spectrum during glucose prediction. For instance, if 1050 nm shows strong glucose-related absorption in a particular individual, its attention score $e_{1050}$ increases, resulting in a higher $\alpha_{1050}$. Because this mechanism is recalculated per input sample, it dynamically adapts to physiological variability such as melanin, hydration, or skin tone. As a result, spectral attention introduces context-aware modulation into the model, improving interpretability and robustness under biological and environmental shifts.
  • Dynamic loss balancing, where $\lambda_{\text{physics}}$ is adaptively scaled based on relative gradient magnitudes to avoid domination by either the data term or the physics term during optimization.
By embedding physics directly into the loss landscape, PINNs reduce the burden on data to learn optical structure from scratch. This allows better generalization under distribution shift—particularly important in field deployment where NIR signals vary due to skin tone, perfusion, and ambient interference.
@@ 3.2.3 Radiative Transfer Equation Models
The Full RTE PINN directly embeds the full radiative transfer equation into its loss function, enabling physically faithful modeling of light propagation in layered, scattering tissue. The governing equation is enforced via a residual minimization loss:
$\mathcal{L}_{\text{RTE}}=\frac{1}{N}\sum_{i=1}^{N}\left\|\mathcal{R}[I^{(i)}]\right\|^{2},$ (30)
where $\mathcal{R}[I]$ is the RTE residual given by:
$\mathcal{R}[I]=\frac{1}{c}\frac{\partial I}{\partial t}+\bm{\Omega}\cdot\nabla I+(\mu_{a}+\mu_{s})I-\mu_{s}\int p(\bm{\Omega},\bm{\Omega}^{\prime})I\,d\bm{\Omega}^{\prime}-S,$ (31)
with $I$ as radiance, $\mu_{a}$ absorption, $\mu_{s}$ scattering, $p(\bm{\Omega},\bm{\Omega}^{\prime})$ as phase function, and $S$ as source term. This formulation captures both directional scattering and spatial diffusion—surpassing the linear Beer–Lambert model and enabling robust generalization in turbid media.
The Selective RTE PINN applies RTE modeling only at critical wavelengths (1050 nm and 1150 nm) where glucose absorption is strongest, reducing computational complexity while maintaining accuracy for glucose-sensitive spectral regions.
@@ 4.1 Overall Performance Comparison
Table 1 presents a comprehensive comparison of all six models evaluated under ultra-realistic synthetic conditions. The Enhanced Beer–Lambert model achieves the best overall performance across multiple metrics, demonstrating the effectiveness of physics-based feature engineering under realistic conditions.
Table 1: Comprehensive Performance Comparison of All Models
@@
The Enhanced Beer–Lambert model demonstrates superior performance with 13.6 mg/dL RMSE, representing a 61% improvement compared to the SDNN baseline. The model achieves 93.8% clinical accuracy (within ±15%) and 95.8% Clarke Error Grid Zone A classification, both exceeding clinical requirements for glucose monitoring devices. Remarkably, this performance is achieved with only 56 parameters and 0.01 ms inference time, making it highly suitable for embedded implementation.
@@
Figure  1 confirms the Enhanced Beer–Lambert model’s dominance across both clinical and computational axes.
Figure 2: Accuracy–complexity Pareto front. X axis: physics depth (0=0\!=\!linear, 1=1\!=\!full RTE); Y axis: performance score (1=1\!=\!best). Bubble area $\propto$ parameter count. The Enhanced Beer–Lambert point (top–left) dominates both axes, illustrating that increased physics depth does not guarantee better real-world accuracy.
The Original PINN achieves competitive performance with 14.6 mg/dL RMSE while incorporating physics constraints. However, the computational complexity increases significantly with 163,000 parameters and 3.8 ms inference time. The Optimized PINN, despite architectural improvements, shows degraded performance under realistic conditions, suggesting that increased model complexity may lead to overfitting when training data contains realistic noise levels.
The RTE-based models (Full and Selective) demonstrate the computational cost of incorporating complete light transport physics. While the Full RTE PINN provides the most physically accurate modeling, its 1.34 million parameters and 15.2 ms inference time make it impractical for embedded deployment. The Selective RTE PINN offers a reasonable compromise with reduced computational requirements while maintaining physics-based modeling for critical wavelengths.
The SDNN baseline shows the poorest performance under realistic conditions, with 35.1 mg/dL RMSE and only 68.2% clinical accuracy. This suggests that purely data-driven approaches without physics constraints are particularly vulnerable to the noise and variations present in realistic implementation scenarios.
For diagnostic and clinical safety analyses, including error grids and agreement metrics, see Appendix A, Figs. 1013.

### v2

# Physics-Informed Neural Networks vs. Physics Models for Non-Invasive Glucose Monitoring: A Comparative Study Under Noise-Stressed Synthetic Conditions

@@ Abstract
Non-invasive glucose monitoring outside controlled settings is dominated by low signal-to-noise ratio (SNR): hardware drift, environmental variation, and physiology suppress the glucose signature in NIR signals. We present a noise-stressed NIR simulator that injects 12-bit ADC quantisation, LED drift, photodiode dark noise, temperature/humidity variation, contact-pressure noise, Fitzpatrick I–VI melanin, and glucose variability to create a low-correlation regime ($\rho_{\text{glucose-NIR}}=0.21$). Using this platform, we benchmark six methods: Enhanced Beer–Lambert (physics-engineered ridge regression), Original PINN, Optimized PINN, RTE-inspired PINN, Selective RTE PINN, and a shallow DNN. The physics-engineered Beer–Lambert model achieves the lowest error (13.6 mg/dL RMSE) with only 56 parameters and 0.01 ms inference, outperforming deeper PINNs and the SDNN baseline under low-SNR conditions. The study reframes the task as noise suppression under weak signal and shows that carefully engineered physics features can outperform higher-capacity models in this regime.
@@ 1 Introduction
Non-invasive glucose monitoring remains a long-standing challenge in optical sensing. Despite decades of research, reliable performance outside controlled settings remains elusive. Near-infrared (NIR) spectroscopy – leveraging glucose’s absorption features in the 850–1150 nm range – is promising, but progress has stalled due to the problem’s intrinsic low SNR.
@@
Recent efforts have turned to Physics-Informed Neural Networks (PINNs), which embed physical constraints (e.g., Beer–Lambert law or full radiative transport equations) into neural architectures. These models aim to improve generalization by aligning learning with known optics. However, their robustness in harsh, noise-stressed settings – and how they compare to simpler models with strong physics-based features – remains poorly understood.
This paper closes that gap by introducing a noise-stressed synthetic data generator that simulates hardware, environmental, and physiological variation. Our simulator includes 12-bit ADC quantization, LED drift, contact-pressure noise, Fitzpatrick I–VI melanin, skin thickness from 0.5–4 mm, and glucose variability. Using this platform, we benchmark six model families:
@@
  • RTE-inspired PINN (physics-regularized light transport)
@@
Our findings reveal that the Enhanced Beer–Lambert model achieves 13.6 mg/dL RMSE with only 56 parameters and 0.01 ms inference time – outperforming deeper PINNs and SDNNs under low-correlation conditions. While our simulator captures a wide spectrum of hardware, environmental, and physiological effects, field deployment may still introduce unmodeled complexities – including sensor-skin misalignment, unmeasured confounders (e.g. sweat, hair, dynamic tissue hydration), and population drift over time.
@@
    We develop a noise-stressed NIR simulator that integrates hardware noise, environmental drift, and physiological diversity – including melanin variation and glucose variability.
@@
    We perform a rigorous side-by-side evaluation of six models under identical low-SNR conditions.
@@
    We show that physics-based feature engineering (Beer–Lambert + PMF) outperforms deeper black-box PINNs under the same noisy conditions.
@@
This work combines Beer–Lambert physics with personalized features (PMFs) in a deployable ridge-regression pipeline for non-invasive glucose sensing.
@@ 2 Background and Related Work
2.1 Near‐Infrared Glucose-Monitoring Fundamentals
@@
These factors reduce raw glucose–NIR correlation from >0.8>\!0.8 in cuvette studies to $\rho\!\approx\!0.2$ in field prototypes, motivating the physics-informed and feature-engineered approaches evaluated in this work.
@@ 2.2 Related Computational Approaches
Deep neural networks (DNN/SDNN). Data-driven models (e.g. Srichan et al. [1]) learn spectral–glucose mappings directly, achieving good results in high-SNR datasets, yet overfit or misgeneralise in noise-stressed regimes.
@@
To date, no study rigorously benchmarks physics-engineered linear models against multiple PINN variants under identical, noise-stressed conditions. The present work fills this gap, showing that carefully engineered Beer–Lambert features coupled with ridge regression can outperform deeper networks in low-SNR settings.
@@ 2.3 Physics-Informed Neural Networks
$\mathcal{L}_{\text{Beer–Lambert}}=\frac{1}{N}\sum_{i=1}^{N}\sum_{\lambda}\bigl|r_{BL}^{(i,\lambda)}\bigr|^{2}.$ (4)
Implementation note. In our baseline PINN, only the Beer–Lambert physics loss in Eq. (4) is enforced.
Radiative-transfer (RTE) PINN (background).
@@
which in full PINN formulations can be enforced via an RTE residual loss. We include this here as theoretical background; our implementation uses a lighter, RTE-inspired regularization described in Section 3.2.3.
Conservation constraints (background).
@@
$\mathcal{L}_{\text{cons}}=\frac{1}{N}\sum_{i=1}^{N}\left|\frac{\partial c^{(i)}}{\partial t}+\nabla\!\cdot\!\bigl(c^{(i)}\mathbf{v}^{(i)}\bigr)\right|^{2},$
where $\mathbf{v}$ is a glucose–transport velocity field. We do not implement this term in the present study.
@@
PINNs excel when the physics is accurate and complete; otherwise the residual term can mis-guide training. Choosing $\lambda_{\text{phys}}$ is non-trivial: too small and the network ignores physics; too large and it overfits imperfect equations. Moreover, solving the full RTE residual requires dense angular quadrature, inflating both parameter count and training cost.
In the glucose-sensing literature, most reported PINNs include only Beer–Lambert penalties and omit scattering or personalised anatomy, leaving open the question of their advantage under low signal-to-noise ratios. Our study quantifies that gap by comparing three PINN variants–including an RTE-inspired PINN–against physics-engineered linear models and an SDNN baseline.
@@ 2.4 Radiative-Transfer Equation Modeling
The Beer–Lambert law assumes a single, straight optical path and no scattering–an assumption violated in biological tissue, where photons scatter tens to hundreds of times before detection. To capture this complexity, the radiative-transfer equation (RTE) models the balance of light energy at every point $\mathbf{r}$, time $t$, and direction $\bm{\Omega}$:
@@
  • - $I(\mathbf{r},\bm{\Omega},t)$ is the radiance – energy per unit area per unit solid angle.
@@
We use RTE-inspired regularization as a soft physics prior in one neural model (3.2.3). This tests whether adding scattering-related structure improves generalisation under noise-stressed optical conditions – and how this compares to empirical compensation in engineered models like EBL.
@@ 3 Methodology
3.1 Noise-Stressed Synthetic Data Generation Framework
Robust deployment demands that algorithms be stress-tested under the same noise, drift, and physiology a wearable will face in field conditions. We therefore build a three-layer simulator—hardware/environment/physiology—and drive plausible glucose trajectories through it. The resulting correlation ceiling ($\rho_{\text{glucose-NIR}}\approx 0.21$) matches values reported for early field prototypes [9, 8].
What we implement. The simulator generates independent snapshot measurements (not continuous time series). Glucose values are sampled from diabetes-status–conditioned distributions with additive meal/stress/time-of-day variation; NIR measurements are then constructed via Beer–Lambert-style absorption plus hardware, environmental, and physiological noise terms. This yields a controlled low-SNR regime for benchmarking model robustness.
@@ 3.1.2 Environmental layer: lab–to–field translation
Context: these ranges cover an ICU ward, a tropical street market, and a Colorado ski slope–three environments where non-invasive monitors have previously failed to generalize reliably.
@@ 3.1.3 Physiology layer: population and state diversity
3.1.4 Glucose snapshots: distributional sampling
@@
Outcome: once all layers interact, raw glucose–NIR correlation collapses from 0.82 (cuvette) to 0.21 (prototype), recreating a low-SNR operating point that has stalled past field deployments.
@@ 3.2.1 Enhanced Beer–Lambert Model
The 56 features include raw absorbances, selected wavelength contrasts, second-order curvature terms, PMF-weighted components, and direct PMF entries–explicitly reflecting all biologically and physically grounded variables used during deployment.
@@ 3.2.2 Physics-Informed Neural Network Architectures
$\mathcal{L}_{\text{PINN}}=\mathcal{L}_{\text{data}}+\lambda_{\text{physics}}\cdot\mathcal{L}_{\text{Beer–Lambert}},$ (24)
@@
$\mathcal{L}_{\text{Beer–Lambert}}=\frac{1}{N}\sum_{i=1}^{N}\sum_{\lambda}\left|\log\left(\frac{I_{0}}{I_{\lambda}^{(i)}}\right)-\varepsilon(\lambda)\cdot\hat{c}_{\text{glucose}}^{(i)}\cdot l^{(i)}\right|^{2}.$ (25)
This acts as a structural prior, penalizing predictions that violate expected absorbance behavior given optical path length $l$ (we set $l=1$ during training, absorbing relative scale into the learned bias). Setting $l=1$ serves as a normalization strategy rather than a strict physical assertion. Since true optical path lengths vary significantly across individuals due to factors like skin thickness, fat distribution, and hydration, fixing $l$ allows us to preserve the linear form of Beer–Lambert while absorbing inter-subject variability into the learned weights. In practice, the physics loss is softly weighted ($\lambda_{\text{physics}}\ll 1$) so that it acts as a regularizer–nudging the model toward optical consistency without overpowering the data-driven objective, while incorporating known extinction coefficients $\varepsilon(\lambda)$.
The Optimized PINN uses a deeper architecture and a multihead attention block to fuse NIR and PMF features, but it retains the same Beer–Lambert physics loss as the baseline. In both cases, the physics term is a fixed-weight regularizer rather than a dynamically balanced loss.
Multihead attention. The optimized model applies multihead attention to the concatenated NIR+PMF embedding, allowing the network to reweight feature subspaces adaptively under low-SNR conditions. This improves feature fusion without altering the underlying Beer–Lambert physics prior.
By embedding Beer–Lambert physics into the loss landscape, PINNs reduce the burden on data to learn optical structure from scratch. This can improve generalization under distribution shift–particularly in low-SNR settings where NIR signals vary due to skin tone, perfusion, and ambient interference.
@@ 3.2.3 Radiative Transfer Equation Models
We implement RTE-inspired models that predict absorption/scattering-related quantities and apply a lightweight physics regularizer rather than a full RTE residual. In practice, the training objective is:
$\mathcal{L}=\mathcal{L}_{\text{data}}+\lambda_{\text{phys}}\mathcal{L}_{\text{reg}},$
where $\mathcal{L}_{\text{reg}}$ encourages stable optical-parameter predictions. This provides a soft physics prior without the computational cost of solving the full RTE residual at every step.
The Selective RTE model applies this physics-inspired regularization primarily to the glucose-sensitive wavelengths (1050 nm and 1150 nm) to reduce computation while maintaining focus on the most informative bands.
@@ 4.1 Overall Performance Comparison
Table 1 presents a comparison of all six models evaluated under noise-stressed synthetic conditions. The Enhanced Beer–Lambert model achieves the best overall performance across error and efficiency metrics, demonstrating the effectiveness of physics-based feature engineering under low-SNR conditions.
Table 1: Performance Comparison of All Models (Error, Efficiency, and Diagnostic Metrics). Clarke A and ±15% are reported for diagnostic comparability.
@@
The Enhanced Beer–Lambert model demonstrates superior performance with 13.6 mg/dL RMSE, representing a 61% improvement compared to the SDNN baseline. This performance is achieved with only 56 parameters and 0.01 ms inference time, making it well-suited to embedded implementation.
@@
Figure  1 confirms the Enhanced Beer–Lambert model’s dominance across both error and computational axes.
Figure 2: Accuracy–complexity Pareto front. X axis: physics depth (0=0\!=\!linear, 1=1\!=\!RTE-inspired); Y axis: performance score (1=1\!=\!best). Bubble area $\propto$ parameter count. The Enhanced Beer–Lambert point (top–left) dominates both axes, illustrating that increased physics depth does not guarantee better field accuracy.
The Original PINN achieves competitive performance with 14.6 mg/dL RMSE while incorporating physics constraints. However, the computational complexity increases significantly with 163,000 parameters and 3.8 ms inference time. The Optimized PINN, despite architectural improvements, shows degraded performance under noise-stressed conditions, suggesting that increased model complexity may lead to overfitting when training data contains noise-stressed levels.
The RTE-based models (RTE-inspired and Selective) demonstrate the computational cost of incorporating light-transport physics. The RTE-inspired PINN provides the most physics-heavy modeling in this study, but its 1.34 million parameters and 15.2 ms inference time make it impractical for embedded deployment. The Selective RTE PINN offers a reasonable compromise with reduced computational requirements while maintaining physics-based modeling for critical wavelengths.
The SDNN baseline shows the poorest performance under noise-stressed conditions, with 35.1 mg/dL RMSE. This suggests that purely data-driven approaches without physics constraints are particularly vulnerable to the noise and variations present in realistic implementation scenarios.
For additional evaluation plots (error grids and agreement analysis), see Appendix A, Figs. 1013.